# AgentOhana: Design Unified Data and Training Pipeline for Effective Agent Learning

### 23 Feb 2024 v1

@@ 2.3 AgentRater
Table 1: Statistics of AgentOhana. AgentOhana consists of data from 10 different environments. #Sampled Trajs indicates the trajectories sampled and filtered from the original environment, #Filtered Trajs indicates the filter trajectories with the AgentRater score $>=4.0$, #Average Turns indicates average number of turns in the filtered trajectories. Among the environments, Operating System, DataBase, Mind2Web and Knowledge Graph are derived from  (Zeng et al., 2023).
@@ 2.4.2 Custom Datasets Creation
class WebshopMultiTurn(AgentModelDatasetBase): …
@@ 3.1 Training
We adopted a supervised fine-tuning approach to enhance the performance of our agent model, xLAM-v0.1, which was initially pre-trained on the Mixtral-8x7B-Instruct-v0.1 model (Jiang et al., 2024). To execute this fine-tuning process, we capitalized on the capabilities of AgentOhana. Our fine-tuning procedure was conducted concurrently on 8 Nvidia H100 GPUs, utilizing the 4-bit quantized QLoRA framework (Dettmers et al., 2023).
@@ 3.2 Benchmarks
In the subsequent sections, we will present experimental evaluations conducted across three benchmarks: Webshop (Yao et al., 2022), HotpotQA (Yang et al., 2018), and MINT-Bench (Wang et al., 2023a). Webshop introduces an online shopping environment that simulates the experience of purchasing specific product items. HotpotQA encompasses multi-hop question-answering tasks, necessitating logical reasoning across multiple Wikipedia passages via the Wikipedia API. We follow the settings established in BOLAA Liu et al. (2023b) for both Webshop and HotpotQA benchmarks. BOLAA’s framework delineates five distinct single-agent settings alongside a multi-agent scenario, providing a structured approach to assess model performance. For the Webshop benchmark, BOLAA comprises 900 user queries, of which we serve 200 as a test subset while the remainder facilitates model development. In the context of HotpotQA, BOLAA samples 300 user questions into three difficulty levels—easy, medium, and hard—with each category containing 100 questions. These questions are exclusively reserved for model testing to ensure a rigorous evaluation process.
We adopt BOLAA’s specified evaluation metrics, namely the average reward, to quantify model performance across benchmarks. Within the Webshop context, the reward metric evaluates the overlapping between attributes of the purchased item and the ground-truth item, thereby measuring model accuracy in identifying user preferences. For HotpotQA, the reward is quantified as the F1 score, reflecting the accuracy of the agent’s predicted answers against the ground-truth responses.
Regarding to the MINT-Bench (Wang et al., 2023a) benchmark, it evaluates LLMs’ ability to solve tasks with multi-turn interactions by using tools and leveraging natural language feedback. The benchmark focuses on reasoning, coding, and decision-making through a diverse set of established evaluation datasets, and carefully curate them into a compact subset for efficient evaluation. The benchmark asks LLMs to solve tasks with different interaction limits from 1 to 5 step and quantify LLMs’ tool-augmented task-solving capability by absolute performance success rate, which measures the percentage of successful task instances as a function of interaction steps.
@@ 3.5 Mint-Bench
Table 4: Testing results on MINT-Bench.

### 26 Feb 2024 v2

@@ 2.3 AgentRater
Table 1: Statistics of AgentOhana. AgentOhana consists of data from 10 different environments. #Sampled Trajs indicates the trajectories sampled and filtered from the original environment, #Filtered Trajs indicates the filter trajectories with the AgentRater score $>=4.0$, #Average Turns indicates average number of turns in the filtered trajectories. Among the environments, Operating System, DataBase, Mind2Web and Knowledge Graph are derived from  (Zeng et al., 2023). Additionally, AgentOhana integrates partial general instruction data sourced from DialogStudio (Zhang et al., 2023), this subset is not represented in the table.
@@ 2.4.2 Custom Datasets Creation
class WebshopMultiTurn(AgentModelDatasetBase):
@@ 3.1 Training
We adopted a supervised fine-tuning approach to enhance the performance of our agent model, xLAM-v0.1, which was initially pre-trained on the Mixtral-8x7B-Instruct-v0.1 model (Jiang et al., 2024). To execute this fine-tuning process, we capitalized on the capabilities of AgentOhana. Our fine-tuning procedure was conducted concurrently on 8 Nvidia H100 GPUs, utilizing the QLoRA framework (Dettmers et al., 2023).
@@ 3.2 Benchmarks
In the subsequent sections, we will present experimental evaluations conducted across four benchmarks: Webshop (Yao et al., 2022), HotpotQA (Yang et al., 2018), ToolEval (Qin et al., 2023), and MINT-Bench (Wang et al., 2023a).
Webshop creates an online shopping environment simulating product purchases, while HotpotQA involves multi-hop question-answering tasks requiring logical reasoning across Wikipedia passages via the Wikipedia API. We adopt BOLAA’s framework Liu et al. (2023b), comprising five single-agent settings and a multi-agent scenario, to evaluate model performance. For the Webshop benchmark, BOLAA comprises 900 user queries, of which we serve 200 as a test subset. For HotpotQA, 300 user questions are sampled into three difficulty levels—easy, medium, and hard—with each category containing 100 questions. These questions are exclusively reserved for model testing to ensure a rigorous evaluation process. We use BOLAA’s evaluation metrics, average reward for Webshop and F1 score for HotpotQA, to measure model performance. In Webshop, the reward metric assesses model accuracy based on the attributes overlapping between the purchased and the ground-truth items, while in HotpotQA, it quantifies the accuracy of agent-predicted answers against ground-truth responses.
ToolEval is designed for real-time assessment of functional calling capabilities via RapidAPI, initially utilizing GPT-3.5-Turbo-16k as its evaluator. However, after careful investigation, we found GPT-3.5-Turbo-16k unreliable for assessing complex function calls and tool usage scenarios. Consequently, we switched to GPT-4-0125-preview as our primary evaluator. Due to the real-time nature of these evaluations, APIs may experience downtime or timeouts, leading to inconsistency in model comparisons across different time frames. To address this, all models are evaluated within the same time frame. Our evaluation employs the default depth-first search-based decision tree methodology, augmented by the Pass Rate metric to assess an LLM’s ability to execute instructions, a fundamental criterion for optimal tool usage. We present our findings at the first level of the ToolEval evaluation, focusing on three distinct scenarios: (1) unseen instructions with the same set of tools, (2) unseen tools within previously seen categories, and (3) unseen tools from entirely new categories that have not been seen before.
MINT-Bench benchmark evaluates LLMs’ ability to solve tasks with multi-turn interactions by using tools and leveraging natural language feedback. The benchmark focuses on reasoning, coding, and decision-making through a diverse set of established evaluation datasets, and carefully curate them into a compact subset for efficient evaluation. The benchmark asks LLMs to solve tasks with different interaction limits from 1 to 5 step and quantify LLMs’ tool-augmented task-solving capability by absolute performance success rate, which measures the percentage of successful task instances as a function of interaction steps.
@@ 3.3 Webshop
3.4 HotpotQA
Unseen Insts & Same Set
Unseen Tools & Seen Cat
Unseen Tools & Unseen Cat
TooLlama V2 (Qin et al., 2023) 0.4385 0.4300 0.4350
GPT-3.5-Turbo-0125 0.5000 0.5150 0.4900
GPT-4-0125-preview 0.5462 0.5450 0.5050
[…]
Table 4: Pass Rate on ToolEval on three distinct scenarios. Bold and Underline results denote the best result and the second best result for each setting, respectively.
3.5 ToolEval
Table 4 displays the results on ToolEval. xLAM-v0.1 surpasses both TooLlama V2 and GPT-3.5-Turbo-0125 across all evaluated scenarios, and outperforms GPT-4-0125-preview in two out of the three settings. This performance indicates xLAM-v0.1’s superior capabilities in function calling and handling complex tool usage tasks. We posit that the model’s performance could be enhanced further through data augmentation involving a variety of prompts.
@@ 3.6 Mint-Bench
Table 5: Testing results on MINT-Bench with different interaction limits from 1 to 5 step.